Assume that A and B are events, with P(A)= 0.6, P (B|A)= 0.2 and P(B|A c)= 0.1. Find the probability for


a. P(B)





1
Expert's answer
2021-06-23T05:39:34-0400
"P(B|A)=\dfrac{P(A\cap B)}{P(A)}"

"=>P(A\cap B)=P(B|A)P(A)"

"P(B|A^C)=\dfrac{P(A^C\cap B)}{P(A^C)}"

"=>P(A^C\cap B)=P(B|A^C)P(A^C)"

"P(A^C)=1-P(A)"

"P(B)=P(A\cap B)+P(A^C\cap B)"

"=P(B|A)P(A)+P(B|A^C)(1-P(A))"


"P(B)=0.2(0.6)+0.1(1-0.6)=0.16"